OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find g(20) + f(6).

OpenStudy (anonymous):

g(20) means put the value 20 in the place of g(x) to get the value

OpenStudy (anonymous):

like g(20)= 3(20)+1=61

OpenStudy (anonymous):

now put f(6) in f(x) and get the value

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

what are u confused in?

OpenStudy (anonymous):

whats the answer suppose to be....

OpenStudy (anonymous):

is that \[f(x)=x^2+1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

final answer should be\[(6)^2+1+61=98\]

OpenStudy (anonymous):

thanks